OpenStudy (anonymous):

If demand is inelastic, the numerical value of elasticity is __________. What do you think? @jim_thompson5910 ? Not sure :(

jimthompson5910 (jim_thompson5910):

go back to the definition of what it means to be inelastic

OpenStudy (anonymous):

okay :) inelastic demand is when suppliers are insensitive to a change in price, so the percentage change in quantity supplied is less than the percentage change in price. so "less than the percentage change in price" is what should go in the blank? @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

well they want a numerical value

OpenStudy (anonymous):

confused:(

jimthompson5910 (jim_thompson5910):

I'm not exactly sure what they are looking for here. The value could range from 0 to 1. hmm

OpenStudy (anonymous):

ohh okay yeah im confused

OpenStudy (anonymous):

thanks though!! if you end up figuring out what it would be, please let me know!! appreciate your help!! :D

OpenStudy (blurbendy):

If demand is inelastic, then the numerical value will be less than 1. Elasticity of demand = ( change in Quantity demanded / The original quanity demand ) * (Price of Original / Change in Price) For example, $1 to $2 4 units to 6 units 2/4 * 1/1 = 2/4 = 1/2 which is less than 1, which means the good is inelastic because the numerical value of elasticity is less than 1

OpenStudy (anonymous):

ahh okay thank you!! @blurbendy :)